OpenStudy (anonymous):

how do you solve csc^2x+2cotx=5

OpenStudy (anonymous):

2cotx = (csc^2x)(sin2x) sin2x = 2sinxcosx csc^2x = 1 / sin^2x 2cotx = 2sinxcosx / sin^2x 2cotx = 2cosx / sinx >>>2cotx

OpenStudy (anonymous):

Thats it

OpenStudy (ranga):

Where did the 5 in the problem go?

OpenStudy (anonymous):

I am not sure how to do it:/ Best of luck.

OpenStudy (anonymous):

\[\csc ^{2}x+2\cot x-5=0\] \[1+\cot ^{2}x+2\cot x-5=0,\cot ^{2}x+2\cot x-4=0\] \[cotx =\frac{ -2\pm \sqrt{4-4*1*-4} }{2*1 }\] \[\cot x=\frac{ -2\pm \sqrt{20} }{2 }=\frac{ -2\pm2\sqrt{5} }{2 }=-1\pm \sqrt{5}\]